### v1

# Globular Clusters GMRT Pulsar Search (GCGPS) II: Discovery of five MSPs in M69 and M70

Abstract
This paper reports recent discoveries from the Globular Clusters GMRT Pulsar Search (GCGPS) survey, which aims to uncover pulsars in the globular clusters (GCs) of the Milky Way using the upgraded Giant Metrewave Radio Telescope (uGMRT). Utilising the Band-4 (550750 MHz) and Band-3 (300500 MHz) receivers, the survey targets GCs accessible to uGMRT (), excluding the declination range that can be covered by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). The survey focuses on GCs that have not previously been searched with comparable sensitivity in these radio frequencies. In this paper, we present the discovery of the five MSPs in two GCs, NGC 6637 (M69) and NGC 6681 (M70), each hosting MSPs identified here for the first time. Observations of M69 led to the discovery of two MSPs: J18313220A (M69A) and J18313220B (M69B), both of which we localize with arcsecond precision using interferometric imaging. Observations of M70 resulted in three new MSPs: J18433217A (M70A), J18433217B (M70B), and J18433217C (M70C). Although direct imaging did not yield precise localizations for these MSPs, we provide initial estimates based on uGMRT beam forming and imaging analysis. Additionally, we present preliminary imaging results for other observed GCs, and in cases of non-detections, we report upper limits on pulsed emission based on the rms noise levels in the image plane.
1 Introduction
The close interaction between an old neutron star (NS) and a low-mass star can lead to the formation of a low-mass X-ray binary (LMXB). In such systems, the NS accretes matter from its companion, gradually spinning up in the process. The high stellar densities and interaction rates found in globular clusters (GCs) (Verbunt_Hut_1987; Pooley_2003; Bahramian_2013) make them ideal environments for the dynamical formation of LMXBs—an occurrence that is rare in the Galactic disk. As a result, GCs host approximately more LMXBs per unit mass than the Galactic disk (Sarazin_2003). Over time, these LMXBs evolve into pulsars with spin periods of just a few milliseconds, known as recycled or millisecond pulsars (MSPs); generally with spin periods ms and low magnetic fields (Becker_1999), usually accompanied by a low-mass white dwarf or a non-degenerate companion (for a review, see Tauris_2023). The abundance of LMXBs in GCs leads to a correspondingly high population of MSPs: of the 700 known MSPs, more than 300 are located in GCs111Based on data from the ATNF pulsar catalogue, https://www.atnf.csiro.au/research/pulsar/psrcat/ (Manchester_2005) and Paulo Freire’s GC pulsar list, see https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html; the total number of known pulsars in GCs is currently 350..
The high stellar densities of GCs mean that some MSP binaries have properties unlike any found in the Galactic disk, such as high orbital eccentricities and massive companions. These binaries result from exchange encounters that happen after the pulsar is recycled. These unusual characteristics make them, potentially, powerful probes for studying gravity in the strong-field regime. For instance, J05144002E in NGC 1851 might be an MSP–black hole system (Barr_2024), which could allow tests of alternative gravity theories beyond the current capabilities of binary pulsar timing (Freire_Wex_2024).
These rare systems highlight the importance of continued pulsar searches in GCs. This is one of the main motivations of the Globular Clusters GMRT Pulsar Search (GCGPS) survey. This survey, including a more complete motivation, target selection and technical details, is discussed in detail by Das_2025, which is Paper I in this series; it is also the designation we will use throughout this work.
In this work (Paper II), we provide an update on the current status of the survey and its data analysis in Section 2. In Section 3, we report the discovery of five MSPs in NGC 6637 (M69) and NGC 6681 (M70). In Section 4, we present the radio imaging analysis of the GCGPS-observed clusters, which helps localize the newly discovered pulsars, and provides a set of unassociated faint radio sources, worth following up for potential new MSPs, along with flux-density upper limits for the non-detections. The final imaging data (in FITS format) associated with this paper are also made publicly available (DOI: 10.5281/zenodo.17781807, private untill published in ApJ), offering a valuable resource for planning more targeted and sensitive follow-up searches. Finally, we summarise our findings in Section 5.
2 OBSERVATIONS AND SURVEY STATUS
The GCGPS observations began in May 2023. In each half-yearly uGMRT cycle, we observe a subset of the selected globular clusters to search for new MSPs. Each GC is observed using two simultaneously formed PA beams, one formed with the central square (CSQ) antennas and the other using antennas up to the third arm, in either Band-3 or Band-4, depending on the target. This configuration ensures the best sensitivity possible with the uGMRT across both the full cluster and its core. For GCs with known MSPs, we record coherently dedispersed (CD) phased array data at median GC DM with a time resolution of 40.96 s and 512 frequency channels across a 200 MHz bandwidth. For GCs with no known MSPs a priori, we record standard phased array(PA) data with 81.92 s time resolution and 4096 channels, without any coherent dedispersion. Alongside beam data, we also record the interformatric data to image the GC field for further analysis. The data analysis procedures followed here are the same as those described in detail in Paper I.
Out of the 31 GCs initially targeted under Phase I (as listed in Table 2 of Paper I), we have so far observed and completed the data analysis of 1 GC in Band-3 and 21 GCs in Band 4. Although Table 1 of Paper I listed NGC 5286 as a Band-3 target, it was observed in Band-4 instead due to feed constraints, and its Band-4 analysis is presented in this paper. Based on this progress, approximately 50% of the planned Phase I survey is complete in Band-3, and about 73% is complete in Band 4. So far, this survey has led to the discovery of seven millisecond pulsars (MSPs) in four globular clusters where no pulsars were previously known222See the survey webpage: http://www.ncra.tifr.res.in/~jroy/GC.html. The first discovery of this survey, J16172258A in NGC 6093, along with the survey details and its timing follow-up, was presented in Paper I. In this paper, we report five additional discoveries in M69 and M70, present their initial study and imaging analysis of all observed and analysed GCs.
Table 1 lists the globular clusters (GCs) observed and analysed until now in this survey (a subset of Phase-I targets), along with their observational details. The remaining Phase I observations, along with the planned Phase II observations—which will include several GCs, which were previously excluded in Phase I due to sensitivity limits or limited sky coverage — are currently ongoing. These Phase II observations are being enabled by the newly commissioned multibeam SPOTLIGHT backend, which offers higher sensitivity aided with multibeam coverage (currently 160 beams), which is much broader than the single-pointing Phase I GCGPS survey (as outlined in Section 3.3 of Paper I). This enhancement in both sensitivity and sky coverage from the Phase I GCGPS observations will allow us to re-observe the TRAPUM targets, along with other targets that we eliminated either due to limited sky coverage or comparatively lower sensitivity and earlier covered ones. The results from these ongoing and upcoming observations will be presented in future publications under this survey.
| GC name | Obs epoch | uGMRT BAND | Obs mode | On-source time | Beam center (RA, Dec) | ||
| (NGC XXXX) | MJD | (BAND-4/3) | (PA/CD) | (sec) | (hh:mm:ss, dd:mm:ss) | () | () |
| NGC 1904 | 60282 | BAND-4 | PA | 9060 | (05 24 11.09, 24 31 29.0) | 112 | 71 |
| NGC 4590 | 60273 | BAND-4 | PA | 8220 | (12 39 27.98, 26 44 38.6) | 413 | 270 |
| NGC 5286 | 60608 | BAND-4 | PA | 6960 | (13 46 26.81, 51 22 27.3) | 361 | 236 |
| NGC 5897 | 60447 | BAND-4 | PA | 6960 | (15 17 24.50, 21 00 37.0) | 201 | 131 |
| NGC 5986 | 60110 | BAND-4 | CD | 4560 | (15 46 03.00, 37 47 11.1) | 173 | 117 |
| NGC 6093 | 60126 | BAND-4 | PA | 8220 | (16 17 02.41, 22 58 33.9) | 132 | 86 |
| NGC 6139 | 60372 | BAND-4 | PA | 7560 | (16 27 40.37, 38 50 55.5) | 1553 | 869 |
| NGC 6293 | 60664 | BAND-4 | PA | 7200 | (17 10 10.20, 26 34 55.5) | 802 | 544 |
| NGC 6333 | 60447 | BAND-4 | PA | 7440 | (17 19 11.26, 18 30 57.4) | 186 | 117 |
| NGC 6355 | 60632 | BAND-4 | PA | 8580 | (17 23 58.59, 26 21 12.3) | 532 | 348 |
| NGC 6388 | 60066 | BAND-4 | PA | 5400 | (17 36 17.23, 44 44 07.8) | 146 | 92 |
| NGC 6528 | 60664 | BAND-4 | PA | 7260 | (18 04 49.64, 30 03 22.6) | 353 | 231 |
| NGC 6541 | 60357 | BAND-4 | PA | 9660 | (18 08 02.36, 43 42 53.6) | 261 | 171 |
| NGC 6553 | 60632 | BAND-4 | PA | 7560 | (18 09 17.60, 25 54 31.3) | 98 | 62 |
| NGC 6569 | 60981 | BAND-4 | PA | 7860 | (18 13 38.80, 31 49 36.8) | 120 | 78 |
| NGC 6637 | 60447 | BAND-4 | PA | 6960 | (18 31 23.10, 32 20 53.1) | 92 | 62 |
| NGC 6638 | 60981 | BAND-4 | PA | 9660 | (18 30 56.10, 25 29 50.9) | 72 | 45 |
| NGC 6642 | 60379 | BAND-4 | PA | 9000 | (18 31 54.10, 23 28 30.7) | 125 | 85 |
| NGC 6652 | 60632 | BAND-4 | CD | 7080 | (18 35 45.63, 32 59 26.6) | 241 | 163 |
| NGC 6681 | 60431 | BAND-4 | PA | 8640 | (18 43 12.76, 32 17 31.6) | 212 | 112 |
| NGC 6723 | 60146 | BAND-4 | PA | 8160 | (18 59 33.15, 36 37 56.1) | 456 | 277 |
| NGC 6809 | 60421 | BAND-3 | PA | 13560 | (19 39 59.71, 30 57 53.1) | 511 | 322 |
3 DISCOVERIES
We now report the discovery of five MSPs in the PA beam data of two GCs with no previously known pulsars and provide a detailed summary of their initial parameters.
3.1 Discoveries in M69
As shown in Table 1, we observed M69 in May 2024 for two hours of on-source time as part of the GCGPS survey. As stated in Section 2, two simultaneous uGMRT Band-4 PA beams were formed (CSQ and the 3rd arm) to observe M69 with optimal sensitivity, aiming to detect new pulsars. We used the YMW16 model predicted DM () and searched over the DM space of 40 to 180 (following the DM plan detailed in Paper I) with a DM step of 0.1 . An initial search on CSQ data led us to the discovery of J18313220A, henceforth M69A. With , this pulsar was detected at a DM value of 82.1 (finer search revealed ), and the discovery SNR from CSQ data was 17.5. After this detection, a finer search over the DM space with a DM step of 0.02 led us to the detection of another candidate, which we later in follow-up observation confirmed to be J18313220B, henceforth M69B. The discovery SNR from the CSQ data for this pulsar ) was 10, and the DM value was 81.92 .
Having 3rd arm beam central theoretical sensitivity times than the CSQ beam central sensitivity, a general search over the DM space with a DM step 0.02 was also performed on the third arm data for pulsars situating close to the GC center, but no promising candidates were detected.
To obtain the initial positional estimates, we folded the 3rd arm data for these two MSPs. The detection significance in both cases was reduced by a factor of 40%60% compared to the CSQ detection, suggesting that the MSPs are located away from the cluster center, where the CSQ beam has 40%60% higher sensitivity than the 3rd arm beam. Based on this estimate, we performed imaging of the field to localize the MSPs; this is discussed in Section 4.1.
During the discovery observation, M69A was detected with a significant acceleration (), confirming that it’s a binary MSP in a relatively compact orbit. Now, for M69B, we could not detect any significant acceleration, indicating that it is an isolated or an MSP in a binary orbit wide enough so that two hours of on-source time is not enough to detect the orbital acceleration. Timing follow-up observations are underway to obtain precise timing solutions for these systems, which we plan to publish in a subsequent paper as part of this survey.
3.2 Discoveries in M70
Similar to the M69 observation, in May 2024 (observation details in Table 1), two simultaneous uGMRT Band-4 PA beams were formed (CSQ and the 3rd arm) to observe M70 with optimal sensitivity, aiming to detect new pulsars. Using the YMW16 model predicted DM (), we searched over the DM space of 30 to 160 with a DM step of 0.1 . The search on the CSQ data led us to discover three MSPs in this GC. PSR J18433217A, henceforth M70A, a MSP having spin period , was discovered at a DM value of with a SNR of . The strongest detection came from J18433217B (M70B, ), at a DM value 71.2 with an SNR of about 34.55. The faintest one is J18433217C (M70C, ), detected in the CSQ data with a SNR of 12.06. After these detections, the more accurate DM for each MSP was determined with PRESTO based search around the detected DM, and presented in Table 2.
All these MSPs were re-detected in the simultaneously recorded 3rd arm data with increased SNR. For all of them, the increment in SNR is about 40%60%, confirming the MSP discovery. This also suggests that they are situated very close to the GC center, and inside the HPBW of the 3rd arm PA beam. Detailed discussion on the localization is done in Section 4.1. Similarly to M69, we also perform a search on the 3rd arm data over with DM step = 0.02 , to detect the MSPs that are very close to the GC center, and might have been missed in the less sensitive CSQ search, but couldn’t detect any promising candidates.
Including the two MSPs in M69, the normalized pulse profiles of all five MSPs are shown in Figure 1, with their corresponding parameters listed in Table 2. In addition, the PREPFOLD outputs for all five MSPs are provided in the Appendix (Section A, Figure 4).
| GC name | Pulsar | Period | DM | Discovery SNR | RA (J2000) and Dec (J2000) | (650 MHz) |
| (NGC XXXX) | (ms) | () | (CSQ, 3rd arm) | (hh:mm:ss, dd:mm:ss) | () | |
| M69/ NGC 6637 | M69A | 3.81 | 82.12 | () | (18:31:22.43, 32.20.47.36) | 224 13 |
| M69B | 4.80 | 81.92 | () | (18:31:23.65, 32:21:10.31) | 278 11 | |
| M70 / NGC 6681 | M70A | 3.93 | 70.65 | () | TBC | TBC |
| M70B | 4.44 | 71.14 | () | TBC | TBC | |
| M70C | 6.06 | 71.06 | () | TBC | TBC |
4 Imaging of the GCGPS observed GCs
From the simultaneously recorded interferometric data, we imaged all the GCGPS observed GCs using the automated pipeline CAPTURE. CAPTURE is designed to take uGMRT interferometric data as input to image them in an automated way when the observation follows a standard imaging observation routine. This pipeline can’t image interferometric data where there is no standard flux calibrator observed. For some of the GCGPS observations, we prioritised having longer on-source time, skipped observing a standard flux calibrator and only observed bright phase calibrators during phasing operations. Now, in order to image these non-standard observations, we modified the pipeline to use the phase calibrator observations for flux calibration by manually providing the calibrator model. This eliminated the need for a standard flux calibration observation and allowed us to perform imaging of the respective GC fields.
4.1 localization of pulsars
As discussed in Section 3.1 and 3.2, both GCs were observed simultaneously using the CSQ and 3rd arm PA beams during the discovery observations. Aided by beam multibeam (CSQ and the 3rd arm) detection, the imaging from the interferometric data allowed us to localize the MSPs within the image field efficiently, provided they were detectable as point sources.
Imaging of M69 revealed multiple radio sources within the field, some of which could correspond to the newly discovered MSPs. To identify the most likely candidates, we used the detection signal-to-noise ratio (SNR) between the CSQ and 3rd arm beams from the discovery observation, along with beam simulation and radiometer equation (Dewey_1985) estimated fluxes. The third arm to CSQ SNR ratios for PSRs M69A and M69B were approximately 0.64 and 0.76, respectively. We simulated the beam responses for both the CSQ and third arm beams during the discovery observation using the uGMRT beam simulation tool uGMRT_beam_fov.py. Based on the observed SNR ratios and the beam response model, we estimated the radial offsets of the MSPs to be within 1530 arcseconds from the cluster center. Guided by these estimates, along with estimated fluxes, imaging revealed two promising point source candidates consistent with the expected positions of the MSPs, along with two additional point sources located slightly farther from the predicted radial range. In subsequent timing follow-up observations, each of the four candidate positions was targeted with two simultaneous beams centerd on their respective locations. The resulting third-arm-to-CSQ SNR ratios confirmed that the two inner candidates correspond to the newly discovered MSPs in M69. The localized positions of these two MSPs and their associated radio continuum sources are shown in Figure 2, and their RA, Dec, and measured flux densities from the images are listed in Table 2. The two outer candidates (marked as “A” and “B” in Figure 2), which were not associated with the discovered MSPs or any other known pulsars, remain strong candidates for potential MSPs.
For the three MSPs in M70 (M70A, M70B, and M70C), the third arm to CSQ detection SNR ratios were approximately 1.39, 1.49, and 1.45, respectively. Similar to our analysis for M69, we simulated the uGMRT beam responses (CSQ and third arm) at the time of the discovery observation for M70 and used the observed SNR ratios to estimate the likely radial positions of the MSPs within the field. The simulations suggest that the expected positions of all three MSPs lie within 10 arcseconds of the cluster center, with the highest probability density concentrated at the very core. Imaging of M70 revealed a relatively strong, extended radio source located near the cluster center, which is unlikely to be associated with any of the MSPs due to its morphology and high flux density. In addition to this, a fainter, compact point source was detected almost exactly at the cluster center, marked as “A” in Figure 3. Given its compactness and alignment with the predicted positions, this source is a promising candidate for any of the three MSPs. No other plausible sources were identified within the cluster region. However, dedicated timing follow-up is required to confirm their actual positions and to establish a definitive association with the imaged source.
4.2 The GC images
We present all twenty-two beam-processed GC images in the Appendix section, grouped into three sets (see Section B, Figures 5, 6, and 7). These images reveal a wealth of information about each GC field, providing a valuable basis for extracting further scientific insights.
The clusters NGC 5986 and NGC 6652 had known MSPs prior to the GCGPS observations. In the case of NGC 5986, the imaging revealed several faint compact sources (with flux densities of a few hundred Jy) located outside the half-light radius, which could be potential MSP candidates. However, the lack of a well-localised position for the previously discovered pulsar in this cluster (J15463747A) did not allow us to associate it with any of the detected sources in the image field. For NGC 6652, similar to the findings of Gautam2022, we detect the source that they associated with J18353259B (marked in the image), but no findings for J18353259A, along with several other faint compact sources that are not associated with any known radio emitters. These additional detections are promising candidates for follow-up studies aimed at identifying new MSPs.
There are now four GCs (NGC 1904, NGC 6093, NGC 6637, and NGC 6681) that host first known MSPs for each GC, discovered through the GCGPS survey. For NGC 1904, the localisation work for the GCGPS-discovered pulsar NGC 1904A (J05242431A) is currently ongoing. We include the image of NGC 1904 in this paper for completeness. In NGC 6093, three MSPs are currently known: the first discovered by the GCGPS survey (PAPER I), and two subsequently found by the TRAPUM survey333See https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. Of these, only NGC 6093A (J16172258A) has a well-localised position (from PAPER I), which we mark in the image. Similarly, for NGC 6637, the localised MSPs discovered in this work are marked in its image. The image of NGC 6681 is also presented, although the positions of the MSPs discovered by TRAPUM in this cluster (three more after the first three discoveries from the GCGPS survey) are not yet publicly available. Across all four clusters, the images reveal several faint yet compact, unassociated radio sources. These may correspond to the GCGPS-discovered MSPs (for NGC 1904 and NGC 6681), or TRAPUM-discovered MSPs (in NGC 6093 and NGC 6681), whose positions are not yet released, or they could represent entirely new, yet-undiscovered MSPs. These sources warrant targeted follow-up observations.
Except for these six GCs, sixteen additional clusters in the GCGPS survey resulted in non-detections and currently do not host any known MSPs. For most of these clusters, the imaging reveals several previously unassociated compact radio sources, which could plausibly include yet-undiscovered MSPs. To enable dedicated follow-up observations of these clusters with deeper sensitivity, we publish the processed FITS images along with this paper (see DOI: 10.5281/zenodo.17781807, private untill published in ApJ ). We appreciate using these images to identify and prioritise promising targets for future searches.
4.3 Non-detection limits
To estimate the non-detection limit of the undiscovered GC MSPs from the GCGPS observations, we utilise imaging data along with the correlation between continuum emission and the pulsar radiometer equation. This approach allows for a more accurate estimation of detection limits via the derived from the GC images as it inherently accounts for sensitivity losses caused by scintillation or the presence of radio frequency interference (RFI), both of which can lead to sub-optimal phasing.
According to Dewey_1985, for continuum emission, the radiometer equation can be written as:
whereas, for time-domain phased-array (PA) beamforming observation, we can write:
Where all the parameters carry their standard meaning. Here, is the number of antennas used for the continuum observation, and is the number of antennas used for the PA/CD beamforming observation. In GCGPS, for continuum recording, we always used all the available antennas at the time of the observation, which typically ranged from 27 to 30. For beamforming observation, for the central square (CSQ) PA/CD beam, the expected number of antennas is 14 (all the central square antennas), and for the third arm (3rd arm) beam, it’s 22 (taking up to the third antenna in each arm for beam formation).
Using Equation 3 and the rms value derived from the imaging data (), we calculate the beam central sensitivity () achieved in our beamformed observations. For each GC, we take into account the number of antennas used for the continuum observations () as well as for the beamforming observations () while calculating the non-detection limits. For the beamforming non-detection limits, we assume a duty cycle () of 0.1 (10%), which is a reasonable assumption for the typical MSP population. In Table 1, for each GC, we present the non-detection limits for both the continuum and the central-square (CSQ) PA/CD beam at the observed band. Similarly, for the third-arm PA/CD beam, we can calculate the achieved beam central sensitivity (non-detection limit) using Equation 3. For the third-arm PA/CD beam, this sensitivity () will scale by a factor of compared to , following the relation , where theoretically and .
5 SUMMARY
This paper presents the discovery of five MSPs from the Globular Clusters GMRT Pulsar Survey (aka GCGPS). All these MSPs are the first MSPs to be discovered in the respective GCs.
The GCGPS observation of the GC M69 resulted in the discovery of two MSPs: M69A is a 3.80 ms MSP (), it was detected with a significant acceleration () during the discovery observation, confirming it’s a binary in a relatively compact orbit; whereas M69B, a 4.80 ms MSP (), was detected with no acceleration, suggesting it is isolated or a binary system in a wide orbit. We localize both MSPs in the image plane (see Figure 2 and Table 2) with arc-sec precision, by simulating the discovery uGMRT beams along with using the simultaneously recorded image data during discovery observation and subsequent follow-up observations.
The observation of M70 led to the discovery of three new millisecond pulsars (MSPs). M70A is a 3.93 ms pulsar with a dispersion measure (DM) of 70.65 pc cm-3. The brightest among them is M70B, a 4.44 ms MSP with a DM of 71.14 pc cm-3, while the faintest is M70C, a 6.06 ms MSP with a DM of 71.06 pc cm-3. None of these sources exhibited detectable acceleration over the 2.5-hour discovery observation, indicating they are either isolated or in binary systems with relatively wide orbits. As in the case of M69, the beam simulation, along with the detection ratios of both beams (CSQ and 3rd arm), indicated that all three MSPs are located within the very central region of M70. Imaging (see Figure 3) revealed a very bright, extended radio source near the center of the globular cluster, which prevented us from confidently associating any of the MSPs with a compact radio source. However, a faint, compact source was also detected almost exactly at the cluster center. This source is a strong candidate for being associated with one of the MSPs. Dedicated timing follow-up is required to determine the precise positions of the MSPs and to confirm any association with the detected radio source.
The imaging of all globular cluster fields observed in the GCGPS survey, including those that did not yield any MSP discoveries (see Figures 5, 6, and 7), revealed several promising targets that could potentially host undiscovered MSPs. These candidates may have been missed in the original GCGPS search due to reduced sensitivity, particularly in the outer regions of the clusters. The processed FITS files corresponding to these images are published with this paper (DOI: 10.5281/zenodo.17781807, private untill published in ApJ) to support more focused, targeted follow-up efforts. Additionally, the imaging data were used to compute non-detection limits for each of the GCGPS-observed globular clusters using the radiometer equation (Equation 3). Based on the respective observational parameters, the non-detection limits (both continuum and beamformed) for each cluster, at the observed uGMRT frequency band, are presented in Table 1.
Appendix A PRESTO PREPFOLD output of all five discovered MSPs
| (a) J18313220A (M69A) | (b) J18313220B (M69B) |
| (c) J18433217A (M70A) | (d) J18433217B (M70B) |
| (e) J18433217C (M70C) |
Appendix B Radio Continuum Images of GCGPS Observed GCs
| (a) NGC1904 | (b) NGC4590 | (c) NGC5286 |
| (d) NGC5897 | (e) NGC5986 | (f) NGC6093 |
| (g) NGC6139 | (h) NGC6293 | (i) NGC6333 |
| (a) NGC6355 | (b) NGC6388 | (c) NGC6528 |
| (d) NGC6541 | (e) NGC6553 | (f) NGC6569 |
| (g) NGC6637 | (h) NGC6638 | (i) NGC6642 |
| (a) NGC6652 | (b) NGC6681 | (c) NGC6723 |
| (d) NGC6809 |